OpenStudy (anonymous):

1 + cos(6x) = ___? a. 2cos^2 (3x) b. 2sin^2 (3x) c. 3sin(2x) d. 3cos(2x)

OpenStudy (nubeer):

cos2x =2cos^2x - 1 2cos^2(x) = 1+cos2x

OpenStudy (anonymous):

sin^2(3x)= 1-cos^2(3x) sin2x= 2sinx cosx cos2x=1- 2sin^2x

OpenStudy (raden):

i just continue what nubber said, hint : 1+cos2x = 2cos^2(x) change x be 3x to both sides :)